A. A football player kicks a 0.43 kg football for a 3-point field goal. If the ball is kicked at 24 m/s at an angle of 53-degrees, how far will it go before landing back on level ground?

Possible answers: a. 24 m b. 28 m c. 34 m d. 45 m e. 55 m

B. If the goal post is 50 yards away (45.72 meters), and its height above the ground is 2.44 m, does his kick make it through the uprights (assuming it does not hit the uprights)?

Possible answers: a. Yes, but just barely clearing the crossbar. b. Yes, by a few meters c. No, it sails under the crossbar. d. Not enough information.

Addendum: Thank you all for your time, I'll be sure to rate!
Answer #1

A) The question is asking about horizontal range which is given as

R = v2 sin 2 / g

R = 242 sin 2 * 53 / 9.8

R = 56.5 m

option (e) is correct

---------------------------------------------------

B) Now, with R = 45.72 m

time in air

t = 45.72 / 24 * cos 53

t = 3.165 sec

Now, we can find vertical displacement of ball

y = vyt - 1/2gt2

y = 24 * sin 53 * 3.165 - 4.9 * 3.1652

y = 11.58 m

this is the height of ball, so

I think option B is correct, the ball will clear the uprights but by few meters.